OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions log base of 4 (x+3) =-2 Help!

OpenStudy (accessdenied):

\[ \begin{split} log_{4}(x+3) &= -2\\ 4^{log_{4}(x+3)} &= 4^{-2}\\ x+3 &= \frac{1}{4^2}\\ x &= \frac{1}{16} - 3\\ x &= \frac{1}{16} - \frac{48}{16}\\ x &= \frac{1 - 48}{16}\\ x &= \frac{-47}{16}\\ \end{split} \] Check: \[ \begin{split} log_{4}(\frac{-47}{16} + 3) = -2\\ log_{4}(\frac{-47}{16} + \frac{48}{16}) = -2\\ log_{4}(\frac{-47 + 48}{16}) = -2\\ log_{4}(\frac{1}{16}) = -2\\ log_{4}(4^{-2}) = -2\\ -2 log_{4}(4) = -2\\ -2 = -2\\ \end{split} \]